# Botender: Supporting Communities in Collaboratively Designing AI Agents through Case-Based Provocations

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The ACM CHI conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Collaborative and social computing systems and tools
Figure 1. Botender is a system that supports users in collaboratively proposing, iterating on, and deploying changes to a community bot powered by LLM-based AI agents. Botender facilitates testing and iterating on the bot’s behavior through algorithmically generated case-based provocations: interaction scenarios designed to spark user reflection and discussion about desirable bot behavior. Users can iterate on the bot and make collective deployment decisions based on these cases.:
@@ 1. Introduction
Automated agents, often referred to as “bots” in online communities, play diverse and important roles across various community platforms (Seering et al., 2018). For example, conversational bots in Discord servers and Slack workspaces are commonly used to increase user engagement by interacting with users, such as sending welcome messages (Hwang et al., 2024; Bali et al., 2023). Moderation bots (Doan and Seering, 2025; Seering et al., 2024), like the Automoderator on Reddit (Jhaver et al., 2019), are widely adopted to enforce subreddit rules by sending warnings and removing posts that violate community standards. Utility-focused bots, such as ClueBot NG on Wikipedia (Geiger and Ribes, 2010; Geiger and Halfaker, 2013), help maintain the quality of community contributions by reverting damaging edits to articles (Kuo et al., 2024; Halfaker et al., 2025; Halfaker and Geiger, 2020). These bots not only perform specific task-oriented functions but also act as social actors in online communities (Seering et al., 2019a), interacting with community members through conversations or other platform actions (Zhang et al., 2020). They are deeply integrated into the sociotechnical infrastructure of online communities and are vital for their growth, maintenance, and flourishing (Seering et al., 2018, 2019a; Geiger and Halfaker, 2013).
@@
Recent advances in Large Language Models (LLMs) have significantly reduced the technical barriers to bot design (Houde et al., 2025; Petridis et al., 2024; Louie et al., 2024; Liu et al., 2025a). This presents an opportunity to support a more community-driven approach, where members are empowered to collectively shape the behavior of bots powered by LLM-based AI agents. [^1] In this paper, we sometimes use “AI agents” and “bots” interchangeably. We generally use “bot” for the user-facing application, and “AI agent” when referring to the underlying implementation of a bot, which may involve one or more AI agents. By viewing bots as shared community infrastructure, such a community-driven approach could help ensure their alignment with a community’s collective values and needs, rather than leaving this in the hands of outsiders or a small group of technical experts. However, supporting communities of non-AI experts in collaboratively designing LLM-based bots for themselves presents several open challenges. First, research shows that non-AI experts often focus narrowly on editing LLM prompts for a single interaction scenario (Subramonyam et al., 2025), failing to test and account for unintended bot behavior across a range of relevant scenarios (Zamfirescu-Pereira et al., 2023b). Moreover, without a coordinated process, differing opinions among community members on how bots should behave can lead to difficulties in achieving consensus and effective collaboration (Fazelpour and Fleisher, 2025; Sorensen et al., 2024). Finally, to support wider community participation, as emphasized in past studies (Kuo et al., 2024), this design process has to be tightly integrated into existing community platforms. These challenges highlight a clear research gap in supporting the participatory design of AI agents within community contexts (Seering et al., 2019a; Long et al., 2017; Hwang et al., 2024).
@@
In this work, we present Botender, a system that enables online communities to collaboratively design and tend to their bots over time. As illustrated in Figure 1, Botender enables users to collaborate on (1) proposing desired changes to the bot’s behavior, (2) iterating on the design of prompt-based bot instructions to operationalize desired changes, through collaborative editing and testing, and (3) deploying changes once users reach some level of consensus on its readiness. Botender supports this iterative, collaborative approach to prompt design through algorithmically generated case-based provocations. These are concrete cases that illustrate how the bot would behave in concrete interaction scenarios, selected to (1) provoke users to reflect on potential gaps between desired bot behavior and the behavior yielded by a given prompt, and (2) reveal potential sources of disagreement about desirable bot behavior within the community. These provocations are designed to support iterative and collaborative prompt design by highlighting interaction scenarios that users may not initially consider. Finally, Botender is designed to be deeply integrated within community platforms to reduce barriers to participation and encourage broader community collaboration in bot design.
@@ 2.1. The Roles of Bots in Online Communities
Bots serve a variety of important roles across different community platforms (Seering et al., 2018). Bots in socially-focused communities, such as Discord servers, usually take on more socially-oriented roles (Seering et al., 2020), while in professional groups like Slack workspaces, they are typically designed to perform task-oriented functions (Ashktorab et al., 2023). Research literature often refers to bots that interact with multiple users simultaneously as multi-party or polyadic bots (Seering et al., 2019a; Zheng et al., 2022). This contrasts with dyadic bots, which primarily engage in one-on-one interactions. Prior research has established taxonomies that categorize the types of content polyadic bots provide to user communities (Long et al., 2017; Seering et al., 2018). For example, Seering et al. classify bot content into five categories (Seering et al., 2018): sharing information (e.g., WikiBot linking to Wikipedia articles [^2] https://www.wikibot.de), sending moderation warnings (e.g., Reddit’s Automoderator [^3] https://www.reddit.com/wiki/automoderator), facilitating user engagement (e.g., MEE6 on Discord welcoming newcomers [^4] https://mee6.xyz), promoting community-approved advertisements (e.g., Nightbot on Twitch advertising streamers’ merchandise [^5] https://nightbot.tv), and running mini-games (e.g., Mudae on Discord for anime character collection [^6] https://mudae.net). Communities typically search for existing bots that roughly meet their needs (Hwang et al., 2024), while HCI researchers have developed a variety of unique bots with specialized roles and functions to address gaps not covered by existing options (Seering et al., 2024; Bali et al., 2023; Zhang and Cranshaw, 2018; Savage et al., 2016; Kim et al., 2020; Liu et al., 2025a).
@@
While online communities generally appreciate the wide range of third-party options available on the market (Hwang et al., 2024), this abundance also creates significant barriers, especially for those without technical backgrounds, when it comes to selecting the “right” bot for their communities (Geiger and Ribes, 2010; Jhaver et al., 2019; Seering et al., 2019b). Since communities often rely on third-party bots created by outsiders who are not part of the community, they frequently end up adopting bots that do not fully align with their specific needs and values, even if these bots provide some customization options (Hwang et al., 2024). As a result, some communities turn to external developers to create custom-tailored bots. However, a study of over two thousand requests on /r/requestabot, a subreddit connecting bot requesters with developers, shows that these external developers still struggle to fully understand the requests due to a lack of community-specific knowledge (Long et al., 2017). In the same vein, without access to the actual community context and community members’ feedback, external developers struggle to iteratively design a bot’s technical functionality and assess its social consequences in communities (Long et al., 2017). While some communities do have members with both the technical skills to build bots and insider knowledge of the community’s needs, the technical barrier to participating in bot design significantly limits broader community participation in shaping the bots’ behavior, which can result in undesirable, community-level consequences that might have been avoided with greater community input (Kraut and Resnick, 2012; Geiger, 2014). For example, bots created by technically inclined Wikipedia patrollers to revert potential vandalism have unintentionally discouraged new contributors, undermining efforts to retain them (Halfaker et al., 2011). Overall, these challenges emphasize the need for research into tools and processes that support more participatory approaches to bot design, allowing communities to collaboratively design bots that better meet their collective needs and values (Seering et al., 2019a; Long et al., 2017; Hwang et al., 2024).
@@ 2.2. Supporting End-Users in Prompt Design
Recent advances in large language models (LLMs) have significantly lowered the technical barriers to designing AI systems capable of more complex tasks (Suresh et al., 2024), such as AI chatbots. This presents an opportunity for communities to design their own LLM-based bots by writing prompts in natural language, without the need for coding. However, crafting effective prompts remains a challenging and unintuitive task for end-users (Zamfirescu-Pereira et al., 2023b). From writing prompts from scratch and iterating on them to assess their downstream impact, prior studies have identified several failure points in prompt design (Zamfirescu-Pereira et al., 2023a). For example, a frequently encountered challenge is that non-AI experts often focus on iterating on their prompts for a specific interaction scenario they have in mind, without considering how these iterations might affect other scenarios (Subramonyam et al., 2025). As a result, prompt iterations can unintentionally worsen outcomes for scenarios previously considered but not revisited, or for relevant scenarios that users had not even considered (Zamfirescu-Pereira et al., 2023b). Insufficient consideration of how a bot might behave across diverse interaction scenarios can lead to prompts that are overly ambiguous, making it difficult for LLMs to distinguish between meaningfully different scenarios (Chen et al., 2024). On the other end of the spectrum, this can also lead to prompts that are worded in a way that is overfit to a specific scenario, so that the LLM is only able to handle a narrow set of interaction scenarios (Subramonyam et al., 2025). Finally, this can lead to prompts that cause unintended downstream consequences in interaction scenarios a user had not considered (Wang et al., 2024b; Zamfirescu-Pereira et al., 2023b).
@@
Building on prior work, Botender aims to support communities in directly collaborating on iterative prompt design through the use of case-based provocations. These provocations are concrete interaction scenarios generated to support reflection on opportunities to improve the bot’s behavior, and to surface differing viewpoints about desirable bot behavior within a community. Most closely related to this concept is a feature in the Gensors system, which automatically generates “edge cases” to help users stress test visual sensing models in situations they might not have considered, to help them to identify unanticipated failure modes (Liu et al., 2025b). In contrast to Gensors’ edge cases, which help individual users debug visual sensors, Botender’s case-based provocations are designed to promote collective reflection on bot behavior across diverse social interaction scenarios—tailored to highlight concrete consequences of known pitfalls in novice prompt design. Botender uses these concrete cases as common ground to facilitate collaborative prompt design. For example, as discussed in Section 4, by voting on cases users can discover where their expectations of bot behavior may differ. The use of cases as common ground in Botender is inspired by their documented success in supporting iterative, collaborative design in other contexts, as discussed in the next subsection.
@@ 2.3. Using Cases to Support Reflection in Collaborative Design Processes
Cases have served as a medium for design and deliberation across many fields (Aamodt and Plaza, 1994). For example, in public policy, concrete cases illustrating how people would be affected by a policy are used as common ground for deliberation and to drive policy changes (Kuo et al., 2025; Epstein et al., 2014; Lehoux et al., 2020; Wright et al., 2020). In the legal context, strategic litigation involves deliberately choosing cases to challenge existing laws, raise public awareness, and promote legal reform (Kirkham, 2023). In HCI research, studies have shown that using concrete scenarios, whether simulated or realistic, helps facilitate the iteration of content moderation rules within online communities (Park et al., 2022; Halfaker et al., 2025; Chen and Zhang, 2023). Across all these contexts, public policy, law, and moderation rules, concrete cases provide a valuable common ground for discussion and for iterating on the high-level, abstract frameworks that govern human behavior.
@@
In the context of LLM-based AI agent design, the framework that governs an agent’s behavior is its prompts. This parallel motivates us to explore how supporting collaborative prompt design can benefit from approaches in other domains, such as public policy, where concrete cases are used to facilitate deliberation and collaborative policy design. For example, PolicyCraft is a system that supports communities in collaboratively proposing, critiquing, and revising regulatory policies through discussion and voting on concrete cases (Kuo et al., 2025). These cases present specific, hypothetical actions by community actors (e.g., community members, businesses, government entities) and allow others to vote on and discuss whether they believe such actions should be permitted in their community. The community then revises its regulatory policies through this collective discussion and consensus. In this work, we explore whether a similar case-based approach can facilitate the collaborative design of AI agent behavior, enabling community members to collectively reflect upon and discuss how they would want a community bot to behave. In contrast to PolicyCraft’s focus on manually-written cases by community members, Botender explores the idea of automatically, dynamically generated case-based provocations (Drosos et al., 2025) that support communities in identifying areas of disagreement and opportunities for bot improvement throughout an iterative, collaborative design process.
@@ 3. Design Goals
Based on a review of prior work highlighting the importance of community-driven approaches to AI agent design, common pitfalls novices encounter when designing agent prompts, and the potential benefits of case-based support, we synthesized the following four design goals for systems that aim to support the collaborative design of AI agents in community settings.
@@
    The system should encourage users to assess the broader impact of their design ideas to support iterative prototyping. Prior research shows that when designing AI agents, non-AI experts often focus narrowly on refining the agent’s behavior in a single scenario, overlooking the broader impact of their prompt designs across a variety of interaction scenarios (Subramonyam et al., 2025; Zamfirescu-Pereira et al., 2023b). To address this, systems should support users in considering how their design may affect a wider range of cases they may not have initially anticipated, to inform design iteration.
@@
    The system should support regression testing to help users prevent the reintroduction of previously resolved issues during iterative agent design. Prior research also shows that non-AI experts often overlook how later design changes can unintentionally reintroduce undesirable agent behaviors they had previously addressed (Zamfirescu-Pereira et al., 2023b). To mitigate this, the system should support regression testing—a practice from software engineering that ensures new codes do not break existing functionality (Elbaum et al., 2000; Rothermel and Harrold, 1997; Elbaum et al., 2014). In the context of collaborative AI agent design, this involves re-running the agent in previously resolved scenarios after making design changes to ensure it still behaves as the community intends.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790500isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Collaborative and social computing systems and tools
Figure 1. Botender is a system that supports users in collaboratively designing community bots. Using Botender, users can propose, iterate on, and deploy changes to a bot powered by LLM-based AI agents. Botender facilitates testing and iterating on the bot’s behavior through algorithmically generated case-based provocations: interaction scenarios designed to spark user reflection and discussion about desirable bot behavior. Users can iterate on the bot and make collective deployment decisions based on these cases.
An illustration of Botender’s user workflow features three main steps: propose, iterate, and deploy, with case-based provocation providing support throughout the entire process.
@@ 1. Introduction
Automated agents, often referred to as “bots” in online communities, play diverse and important roles across various community platforms (Seering et al., 2018). For example, conversational bots in Discord servers and Slack workspaces are commonly used to increase user engagement, such as by sending welcome messages (Hwang et al., 2024; Bali et al., 2023). Moderation bots (Doan and Seering, 2025; Seering et al., 2024), like the Automoderator on Reddit (Jhaver et al., 2019), are widely adopted to enforce subreddit rules by sending warnings and removing posts that violate community standards. Utility-focused bots, such as ClueBot NG on Wikipedia (Geiger and Ribes, 2010; Geiger and Halfaker, 2013), help maintain the quality of community contributions by reverting damaging edits to articles (Kuo et al., 2024; Halfaker et al., 2025; Halfaker and Geiger, 2020). These bots not only perform specific task-oriented functions but also act as social actors in online communities (Seering et al., 2019a), interacting with community members through conversations or other platform actions (Zhang et al., 2020). They are deeply integrated into the sociotechnical infrastructure of online communities and are vital for their growth, maintenance, and flourishing (Seering et al., 2018, 2019a; Geiger and Halfaker, 2013).
@@
Recent advances in Large Language Models (LLMs) have significantly reduced the technical barriers to bot design (Houde et al., 2025; Petridis et al., 2024; Louie et al., 2024; Liu et al., 2025b). This presents an opportunity to support a more community-driven approach, where members are empowered to collectively shape the behavior of bots powered by LLM-based AI agents. [^1] In this paper, we sometimes use “AI agents” and “bots” interchangeably. We generally use “bot” for the user-facing application, and “AI agent” when referring to the underlying implementation of a bot, which may involve one or more AI agents. By viewing bots as shared community infrastructure, such a community-driven approach could help ensure their alignment with a community’s collective values and needs, rather than leaving this in the hands of outsiders or a small group of technical experts. However, supporting communities of non-AI experts in collaboratively designing LLM-based bots for themselves presents several open challenges. First, research shows that non-AI experts often focus narrowly on editing LLM prompts for a single interaction scenario (Subramonyam et al., 2025), failing to test and account for unintended bot behavior across a range of relevant scenarios (Zamfirescu-Pereira et al., 2023b). Moreover, without a coordinated process, differing opinions among community members on how bots should behave can lead to difficulties in achieving consensus and effective collaboration (Fazelpour and Fleisher, 2025; Sorensen et al., 2024). Finally, to support wider community participation, as emphasized in past studies (Kuo et al., 2024), this design process has to be tightly integrated into existing community platforms. These challenges highlight a clear research gap in supporting the participatory design of AI agents within community contexts (Seering et al., 2019a; Long et al., 2017; Hwang et al., 2024).
@@
In this work, we present Botender, a system that enables online communities to collaboratively design and tend to their bots over time. As illustrated in Figure 1, Botender enables users to collaborate on (1) proposing desired changes to the bot’s behavior, (2) iterating on the design of prompt-based bot instructions to operationalize desired changes, through collaborative editing and testing, and (3) deploying changes once users reach some level of consensus on its readiness. Botender supports this iterative, collaborative approach to prompt design through algorithmically generated case-based provocations. These are concrete cases that illustrate how the bot would behave in concrete interaction scenarios, selected to probe the boundaries of users’ intents by provoking user reflection on (1) potential gaps between users’ desired bot behavior and the behavior yielded by a given prompt, and (2) potential sources of disagreement about desirable bot behavior within a community. These provocations are designed to support iterative and collaborative prompt design by highlighting a given prompt’s behavioral consequences in interaction scenarios that users may not initially consider. Finally, Botender is designed to be deeply integrated within community platforms to reduce barriers to participation and encourage broader community collaboration in bot design.
@@ 2.1. The Roles of Bots in Online Communities
Bots serve a variety of important roles across different community platforms (Seering et al., 2018). Bots in socially-focused communities, such as Discord servers, usually take on more socially-oriented roles (Seering et al., 2020), while in professional groups like Slack workspaces, they are typically designed to perform task-oriented functions (Ashktorab et al., 2023). Research literature often refers to bots that interact with multiple users simultaneously as multi-party or polyadic bots (Seering et al., 2019a; Zheng et al., 2022). This contrasts with dyadic bots, which primarily engage in one-on-one interactions. Prior research has established taxonomies that categorize the types of content polyadic bots provide to user communities (Long et al., 2017; Seering et al., 2018). For example, Seering et al. classify bot content into five categories (Seering et al., 2018): sharing information [^2] E.g., WikiBot on Discord links Wikipedia articles to messages: https://www.wikibot.de, sending moderation warnings [^3] E.g., Automoderator on Reddit sends moderation warnings to users (Jhaver et al., 2019), facilitating user engagement [^4] E.g., MEE6 on Discord welcomes newcomers to the server (Lee et al., 2025), promoting community-approved advertisements [^5] E.g., Nightbot on Twitch promotes advertisements for streamers’ merchandise (Seering et al., 2020), and running mini-games [^6] E.g., Mudae on Discord facilitates anime character collection games (Hwang et al., 2024). Communities typically search for existing bots that roughly meet their needs (Hwang et al., 2024), while HCI researchers have developed a variety of unique bots with specialized roles and functions to address gaps not covered by existing options (Seering et al., 2024; Bali et al., 2023; Zhang and Cranshaw, 2018; Savage et al., 2016; Kim et al., 2020; Liu et al., 2025b).
@@
Although many bots are available on the market, communities often struggle to find the “right” bot for their needs and frequently end up adopting bots that misalign with their specific requirements (Geiger and Ribes, 2010; Jhaver et al., 2019; Seering et al., 2019b). This misalignment occurs because existing third-party bots typically offer limited customization options, and developers often lack the community-specific knowledge and capacity to address tailored requests over time (Hwang et al., 2024). For example, a study of more than two thousand requests on /r/requestabot, a subreddit that connects bot requesters with developers, found that external developers who lack community-specific knowledge often struggle to fully understand requests and are unable to develop bots that are tailored to the needs of individual communities (Long et al., 2017). While some communities do have members with both the technical skills to build bots and insider knowledge of the community’s needs, the technical barrier to participating in bot design significantly limits broader community participation in shaping the bots’ behavior, which can result in undesirable, community-level consequences that might have been avoided with greater community input (Kraut and Resnick, 2012; Geiger, 2014). For example, bots created by technically inclined Wikipedia patrollers to revert potential vandalism have unintentionally discouraged new contributors, undermining efforts to retain them (Halfaker et al., 2011). Overall, these challenges emphasize the need for research into tools and processes that support more participatory approaches to bot design, allowing communities to collaboratively design bots that better meet their collective needs and values (Seering et al., 2019a; Long et al., 2017; Hwang et al., 2024).
@@ 2.2. Supporting End-Users in Prompt Design
Recent advances in LLMs have significantly lowered the technical barriers to designing AI systems capable of more complex tasks (Suresh et al., 2024), such as AI chatbots. This presents an opportunity for communities to design their own LLM-based bots by writing prompts in natural language, without the need for coding. However, crafting effective prompts remains a challenging and unintuitive task for end-users (Zamfirescu-Pereira et al., 2023b). From writing prompts from scratch and iterating on them to assess their downstream impact, prior studies have identified several failure points in prompt design (Zamfirescu-Pereira et al., 2023a). For example, a frequently encountered challenge is that non-AI experts often focus on iterating on their prompts for a specific interaction scenario they have in mind, without considering how these iterations might affect other scenarios (Subramonyam et al., 2025). As a result, prompt iterations can unintentionally worsen outcomes for scenarios previously considered but not revisited, or for relevant scenarios that users had not even considered (Zamfirescu-Pereira et al., 2023b). Insufficient consideration of how a bot might behave across diverse interaction scenarios can lead to prompts that are overly ambiguous, making it difficult for LLMs to distinguish between meaningfully different scenarios (Chen et al., 2024). On the other end of the spectrum, this can also lead to prompts that are worded in a way that is overfit to a specific scenario, so that the LLM is only able to handle a narrow set of interaction scenarios (Subramonyam et al., 2025). Finally, this can lead to prompts that cause unintended downstream consequences in interaction scenarios a user had not considered (Wang et al., 2024b; Zamfirescu-Pereira et al., 2023b).
@@
Building on prior work, Botender aims to support communities in directly collaborating on iterative prompt design through the use of case-based provocations. These provocations are concrete interaction scenarios generated to support reflection on opportunities to improve the bot’s behavior, and to surface differing viewpoints about desirable bot behavior within a community. Most closely related to this concept is a feature in the Gensors system, which automatically generates “edge cases” to help users stress test visual sensing models in situations they might not have considered, to help them to identify unanticipated failure modes (Liu et al., 2025a). In contrast to Gensors’ edge cases, which help individual users debug visual sensors, Botender’s case-based provocations are designed to promote collective reflection on bot behavior across diverse social interaction scenarios—tailored to highlight concrete consequences of known pitfalls in novice prompt design. Botender uses these concrete cases as common ground to facilitate collaborative prompt design. For example, as discussed in Section 4, by voting on cases users can discover where their expectations of bot behavior may differ. The use of cases as common ground in Botender is inspired by their documented success in supporting iterative, collaborative design in other contexts, as discussed in the next subsection.
@@ 2.3. Using Cases to Support Reflection in Collaborative Design Processes
Cases have served as a medium for design and deliberation across many fields (Aamodt and Plaza, 1994; Epstein et al., 2014; Lehoux et al., 2020; Wright et al., 2020; Kuo et al., 2023), including HCI research. Building on prior HCI scholarship, we understand cases as concrete scenarios within specific problem situations that make abstract concepts tangible for reasoning, communication, and negotiation (Chen and Zhang, 2025; Kuo et al., 2025). For example, to establish common ground for public policy deliberation, HCI research has used concrete cases to illustrate how individuals might be affected by such policies, which are typically more abstract and difficult to reason about compared to specific cases (Kuo et al., 2025). HCI researchers have also brought strategically-selected legal cases that are more accessible to the general public before the judicial system to challenge existing laws and advocate for legal reform (Kirkham, 2023; Yang et al., 2024). Similarly, in online communities such as Wikipedia and Reddit, reflecting on concrete cases of user content enables moderators to refine moderation rules to account for scenarios they had not initially considered when creating these rules (Park et al., 2022; Halfaker et al., 2025; Kuo et al., 2024; Fan and Zhang, 2020; Koshy et al., 2025; Chen and Zhang, 2023). Across these contexts—public policy, law, and moderation rules—concrete cases provide a valuable common ground for reflection and discussion, to inform policy iteration. This broader idea of iteratively developing general principles based on judgments about concrete cases aligns with the method of reflective equilibrium, proposed by political philosopher John Rawls (Knight, 2017). In this method, the cases most useful for discussing and refining principles are often not those that conform to existing principles, but rather those where case-level judgments conflict with current principles, or cases that elicit divergent judgments among discussants (Kuo et al., 2025). Such cases push the group to reflect on their perspectives, and iteratively refine high-level principles and/or case-level judgments until they are in alignment (or equilibrium) (Chen and Zhang, 2025, 2023).
@@
In the context of LLM-based AI agent design, the principles that govern an agent’s behavior are commonly expressed as prompts. This parallel motivates us to explore how supporting collaborative prompt design can benefit from approaches in other domains, such as public policy, where concrete cases are used to facilitate deliberation and collaborative policy design. For example, PolicyCraft is a system that supports communities in collaboratively proposing, critiquing, and revising regulatory policies through discussion and voting on concrete cases (Kuo et al., 2025). These cases present specific, hypothetical actions by community actors (e.g., community members, businesses, government entities) and allow others to vote on and discuss whether they believe such actions should be permitted in their community. The community then revises its regulatory policies through this collective discussion and consensus. In this work, we explore whether a similar case-based approach can facilitate the collaborative design of AI agent behavior, enabling community members to collectively reflect upon and discuss how they would want a community bot to behave. In contrast to PolicyCraft’s focus on manually-written cases by community members, Botender explores the idea of automatically, dynamically generated case-based provocations that support communities in identifying areas of disagreement and opportunities for bot improvement throughout an iterative, collaborative design process.
@@ 3. Design Goals
Based on our review of prior work, we synthesized the following four design goals for systems that aim to support the collaborative design of AI agents in community settings.
@@
    The system should encourage users to assess the broader impact of their design ideas to support iterative prototyping. As discussed in Section 2.2, when designing AI agents, non-AI experts often focus narrowly on refining the agent’s behavior in a single interaction scenario, overlooking the broader impact of their prompt designs across a variety of interaction scenarios (Subramonyam et al., 2025). As a result, prompt iterations can unintentionally worsen outcomes for other relevant scenarios that users had failed to consider (Zamfirescu-Pereira et al., 2023b). To address this, systems should support users in considering how their design may affect a wider range of cases they may not have initially anticipated, to inform design iteration.
@@
    The system should support regression testing to help users prevent the reintroduction of previously resolved issues during iterative agent design. As discussed in Section 2.2, non-AI experts often focus on refining an agent’s behavior in the interaction scenario they are currently considering, overlooking how their refinements may impact interaction scenarios they had previously considered (Subramonyam et al., 2025; Zamfirescu-Pereira et al., 2023b). As a result, prompt iterations may unintentionally reintroduce undesirable agent behaviors in previously addressed scenarios. To mitigate this issue, prior HCI research (Zamfirescu-Pereira et al., 2023b) suggests that systems should support regression testing—a practice from software engineering to ensure that new code does not break existing functionality (Elbaum et al., 2000; Rothermel and Harrold, 1997; Elbaum et al., 2014). In the context of collaborative AI agent design, this involves re-running the agent in previously resolved scenarios after making design changes to ensure it still behaves as the community intends.